Report compiled by David Watters (Programme Manager) for Alex Carter (Programme Sponsor)
Section 1: Programme Summary
Programme Manager's Commentary
There has been progress on the majority of COM projects this month, but there is still a feeling that circumstances are leading to a sequence of small delays that are compounding to create a sense of frustration around some of the projects. This is despite the fact that work is still getting done and the various project teams are chipping away at the different tasks.
The programme has, again, more Amber than Green projects, with the current reasons for these being addressed. The Programme and Portfolio Managers are continuing to ensure that things do not become stuck and drift into the next financial year. The Qlikview project is still planned for early June and a project request will be submitted soon.
One final point is that the Programme budget has now been fully utilised (with a transfer to COM045) and currently has a small deficit - this is under discussion with senior management.
The IDM DB migration project has resolved the issue encountered in integration testing and the project team are now advancing towards UAT. This has led to some modifications around the final delivery date and the effort required.
The O365 Security Hardening project has had to be adjusted dates-wise again. This is because of the final checks and tasks needing to be extended. The Amber status is due to these modifications.
The Staffmail migration work is back to Green. There are plans to complete the last migrations and draw the project to a close over the next month or so.
The Wiki Crowd Implementation project has been reported as Amber because of the issue around merging users still waiting for input towards resolution from Atlassian. This, and subsequent leave in the project team, has meant that there will be a need to shift the previously published milestones.
The VRS Rewrite project is still AMBER due to the delay on the design decision on the VRS components. The project may return to GREEN next month if this is settled. There continues to be a lot of effort put into advancing this one.
New project milestones agreed by project team and WIS. Budget review has taken place and Programme Manager is in discussion with PMO, Portfolio Manager and Head of Service Management on source of extra days
Integration testing caught and fixed a SOA issue that prevented new students receiving notifications after the database upgrade.
Integration testing signed off. Project has moved to the UAT phase
Budget and timeline review has taken place. Milestone changes were agreed with the project team to extend the Integration, Deployment and Closure milestones of the project by just over one month. This was due to the issues found during integration testing which have now been resolved.
Meeting held 26 April to discuss changes to auto-forwarding
Agreed that managing the groups would be done under the following:
All students exempt
The expectation had been to change the auto-forwarding default allow setting so it matched the Microsoft setting. What hadn’t been intended was to change the groups that had set up for students and staff who were already forwarding. Plus it had been found that some staff who were forwarding no longer forward, staff have left, new staff who are forwarding but not in our list Therefore sponsor has asked for a review of the groups and a refresh of their membership
Update membership of groups
Get new list of who actually does have a forward set and agree it with Alex
Check Johns list to see how many people in support groups versus schools
Set up a functional account for comms
Set up a SharePoint site for email service
Decide a deployment date
Service Management comms (Schools and Colleges Leads - copying in Abdul Majothi, Relationship manager; Users in the current email forwarding group) to be sent out 2weeks before deployment date
Project comms (Helpline and EUSA) to be sent out 2weeks before deployment date
Commentary: Project is AMBER
Description: Additional work to get auto-forwarding delivered
Impact: Unlikely to meet the milestones with the work that has now to be done to the groups and set up of accounts. Budget is on the edge of the 10% allowance
Plan:New milestones will be submitted. PM will monitor costs and submit a piccl if likely to go over
All, bar one, Informatics accounts have been successfully migrated and signed off.
Physics accounts have been scheduled for 4th, 11th and 18th May 2021
Clear down off all accounts has been completed
Historical migration of first batch has been completed
Historical migration or remaining two batches is scheduled
Work has progressed with reducing the number of remaining accounts on the Staffmail Service. A final series of migrations will happen at the end of May/beginning of June to clear the stragglers. The majority of these accounts are functionals.
There are no issues to report.
Complete planned Physics migrations
Continue to reduce the number of accounts remaining
Start migrating remaining accounts.
After rescheduling the remaining Physics migrations, the project RAG is now GREEN
Atlassian's input is required into the merge of users between Central Auth, Wiki and Crowd. We have 308k Wiki users and 78k users from Central Auth and these need to be synched. Due to this being a DEV environment Atlassian have been slower in providing updates as their priorities are with customers with live issues. Dev Tech require Atlassian's input in verifying if SQL code written to merge will work.
Dev Tech resource is on annual leave until 17/05. Milestones will need to be replanned with DEV signoff to be the 21/05
Due to the issues with the merge, the project team have agreed that the deployment to LIVE will now need to take place in June
Run the merge code for DEV
Milestone updates to be agreed at WIS. Indicative milestones are below
Start the Crowd implementation to TEST
Amber due to milestone revisions required. Merge of users between Wiki / Central Auth into Crowd requires Atlassian's input into the SQL written to do this, however this being the DEV environment means a lower priority for Atlassian, and a longer turnaround time to answer questions. Dev Tech resource has annual leave from 05/05 to 16/05.
Work on developing the IDM functional areas has continued.
Work on the Design document for VRS has been delayed due to annual leave. However, the draft VRS Design document has been shared with the Head of Development Services, who will provide an estimate for the VRS build using Python instead of O365 tools.
A second round of comms have been issued with more specific questions around service usage and allocation. The project will schedule workshops with key stakeholders when this information has been collated.
Requirements gathering for bulk upload has started.
Again this month, the decision regarding the VRS solution is taking longer than expected.
The technical resource assigned to the project could not honour the bookings. Resources bookings have been re-requested and another colleague from DevTech assigned to the project.
Complete and estimate the design document for VRS components
make a decision on the solution for the VRS components
start stakeholder workshops
Continue building the IDM components
Continue with technical build documentation
The project is AMBER due to the continued delay on the design decision on the VRS components.